Situation: Brazil and Chile can produce cloth and wheat. In Chile, the labor requirement for a yard of cloth is 6 hours while the requirement for a bushel of wheat is 9 hours. In Brazil, the requirement for a yard of cloth is 6 hours while that for a bushel of wheat is 10 hours.

Question: Is it possible for the wheat-exporting country to charge 12 bushels/yard for its wheat? Explain briefly.
Answer #1

Opportunity cost of 1 bushel of wheat in Brazil = Labor requirement for wheat/Labor requirement for cloth = 10/6 = 1.67 yard of cloth

Opportunity cost of 1 bushel of wheat in Chile = Labor requirement for wheat/Labor requirement for cloth = 9/6 = 1.5 yard of cloth

So, Chile will export wheat because it has lower opportunity cost of wheat and thus it has a comparative advantage in it.
Trade ratio must lie between 1.67 yard of cloth per bushel and 1.5 yard of cloth per bushel, or (1/1.67) = 0.6 bushel/yard and (1/1.5) = 0.67 bushel/yard. Thus,  12 bushels/yard does not lie within the trade ratio. So, it is not possible for the wheat-exporting country to charge 12 bushels/yard for its wheat.